OpenStudy (anonymous):

On a map, 1 in. represents 20 mi. Two cities are in. apart on the map. How far apart are the actual cities? A.23 1/2 mi B.60 1/2mi C.70 mi D.75mi please help!!

OpenStudy (jdoe0001):

"Two cities are in. apart on the map." ^ ?

OpenStudy (anonymous):

what is the missing thing ? its hard to solve with no given

OpenStudy (anonymous):

3and 1/2 sorry

OpenStudy (anonymous):

sooo do any of you know?

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} \textit{map inches}&\textit{actual miles}\\ 1&20\\ x&3\frac{1}{2} \end{array}\implies \cfrac{1}{x}=\cfrac{20}{3\frac{1}{2}}\)

OpenStudy (jdoe0001):

solve for "x"

OpenStudy (anonymous):

i do not know the answer?!

OpenStudy (jdoe0001):

hmmmm shoot I got the wrong... ... one sec... the wrong item in the wrong column =)

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} \textit{map inches}&\textit{actual miles}\\ 1&20\\ 3\frac{1}{2}&x \end{array}\implies \cfrac{1}{3\frac{1}{2}}=\cfrac{20}{x}\)

OpenStudy (jdoe0001):

now just solve for "x", tis a plain linear equation simplification

OpenStudy (anonymous):

we cannot give out the answer however we can help you get it and understand it @jdoe0001 gave you a great way to start and a good format its not so hard man

OpenStudy (anonymous):

the answer is 70 correct

OpenStudy (jdoe0001):

how did you get 70?

OpenStudy (anonymous):

i am a girl!!!

OpenStudy (jdoe0001):

\(\Large ?\)

OpenStudy (anonymous):

girl*

OpenStudy (anonymous):

i did 3 1/2 broke it into a decimal and then i multiplyed 3.5 times 20 easy and it is correct right?

OpenStudy (anonymous):

bye and thank you guys soo much!!

OpenStudy (jdoe0001):

hehehyes,it's 70 =)